### v1

# SentenceVAE: Faster, Longer and More Accurate Inference with Next-sentence Prediction for Large Language Models

Abstract
Contemporary large language models (LLMs) predominantly utilize a next-token prediction method for inference, which significantly impedes their processing speed. In this paper, we introduce a novel inference methodology termed next-sentence prediction, aimed at enhancing the inference efficiency of LLMs. We present SentenceVAE, a tiny model consisting of an encoder and a decoder. The encoder effectively condenses the information within a sentence into a singular token, while the decoder reconstructs this compressed data back into its original sentential form. By integrating SentenceVAE into the input and output layers of LLMs, we develop Sentence-level LLMs (SLLMs) that employ a sentence-by-sentence inference approach, markedly accelerating inference speeds. SentenceVAE also maintains the integrity of the original semantic content by segmenting the text into sentences, thereby preserving accuracy while boosting inference speeds. Compared to traditional LLMs, SLLMs process fewer tokens over equivalent context lengths, significantly reducing memory demands for Self-Attention computations and facilitating the handling of longer contexts. Our experimental findings reveal that this method can increase inference speeds by 204$\sim$365%, reduce perplexity (PPL) to 46$\sim$75of its original metric, and decrease memory overhead by 86$\sim$91for the same context length. The advantages of this approach are further amplified with increases in model parameters.
Introduction
Figure 1: The schematic form of SentenceVAE. It clearly illustrates that the encoder of SentenceVAE can compress the information contained within a sentence into a single token, and the decoder can restore the compressed token back to its original sentence form.
Large language models (LLMs) have demonstrated remarkable proficiency in understanding human intent through the acquisition of knowledge from vast and diverse datasets. These models provide rich and accurate information and responses, making them indispensable in applications such as machine translation (Zhang, Haddow, and Birch 2023; Wang et al. 2023), chatbots (Islam and Moushi 2024; Wang et al. 2024), and question-answering systems (Singhal et al. 2023; Lazaridou et al. 2022). The predominant approach to inference in LLMs involves next-token prediction, where the model generates tokens sequentially. While this token-by-token generation method has achieved significant performance, it is inherently limited in that it produces only one token per inference step. This limitation results in substantial computational overhead and increases the overall inference time, thereby constraining the scalability and efficiency of LLMs. As a result, there is a pressing need to explore alternative methods that can mitigate these limitations while maintaining or improving upon the accuracy and responsiveness of LLMs.
To enhance the inference efficiency of Large Language Models (LLMs) and reduce computational costs, Gloeckle et al. (Gloeckle et al. 2024) proposed training LLMs to predict multiple tokens simultaneously, aiming to accelerate the inference process. However, the number of tokens predicted at once is fixed. For instance, when this number is set to five, regardless of the input text, it is forcibly segmented into units of five tokens. While this multi-token prediction method improves the inference speed of LLMs, it may disrupt the original semantics of the text by aggregating unrelated tokens together, potentially compromising inference accuracy.
To enhance inference speed while maintaining accuracy, we explore methods for adaptively selecting the optimal number of predicted tokens based on the semantic content of the text. Consequently, we introduce SentenceVAE, which comprises a sentence encoder and a sentence decoder. The schematic of SentenceVAE is illustrated in Fig. 1. The sentence encoder is designed to condense the information of an entire sentence into a single token, while the sentence decoder reconstructs this compressed token back into a sentence. By integrating the encoder and decoder of SentenceVAE into the input and output layers of Large Language Models (LLMs), these models can perform next-sentence prediction, thereby accelerating the reasoning process. Moreover, since SentenceVAE segments text at the sentence level, the semantic integrity is better preserved, ensuring the accuracy of the inferrence.
Specifically, prior to inputting text into LLMs, SentenceVAE segments the text into sentences and utilizes the sentence encoder to compress the information contained within each sentence into a single token. These compressed tokens are then fed as input to the LLMs. The LLMs predict their output based on these compressed tokens. Finally, the sentence decoder decodes the predicted token into the final output of the LLMs.
The contribution of the proposed method can be described as follows:
  • We propose SentenceVAE, a model capable of compressing the information content of a sentence into a single token, and conversely, expanding such tokens to reconstruct sentences with high informational fidelity.
  • We embed the encoder and decoder components of SentenceVAE into the input and output layers of LLMs, enabling these models to implement an effective inference method for next-sentence prediction.
  • Extensive experiments conducted on various LLMs have demonstrated that the proposed method significantly accelerates inference speed while preserving the accuracy of these models.
Figure 2: (a) The schematic form of traditional LLM. (b) The schematic form of Setence-level LLM (SLLM), which embedded with SentenceVAE. It can be clearly seen that, unlike the next-token inference method employed by traditional large language models (LLMs), the proposed method adopts a next-sentence prediction approach, which significantly reduces the number of inference iterations and the overall inference cost of LLMs.
Related Work
Large language models: In recent years, large language models (LLMs) have achieved remarkable performance in the field of natural language processing (NLP). However, these models typically possess a vast number of parameters from billions to hundreds of billions, exemplified by models such as TeleChat (1B-52B) from TeleAI (Wang et al. 2024; Li et al. 2024b), Llama (7B-405B) from Meta (Touvron et al. 2023a, b), InternLM (1.8B-20B) from Shanghai AI Lab (Cai et al. 2024), etc. The sheer magnitude of these parameters renders the inference process computationally intensive and time-consuming. Moreover, most LLMs employ a one-token prediction method, where the model predicts only one token at a time. This token-by-token approach further exacerbates the inference time of LLMs. As a result, accelerating the inference speed of LLMs while maintaining their accuracy presents a significant challenge for researchers in the field.
Multi token prediction: The transformer model enables the parallel processing of sequences through its self-attention mechanism, providing a technical foundation for the implementation of multi-token prediction (Vaswani et al. 2017). The BERT model utilizes bidirectional encoder representations to capture contextual information, further solidifying the basis for the implementation of multi-token prediction (Devlin et al. 2018). The T5 model incorporates multi-token prediction during the training process for the first time, significantly enhancing the coherence and quality of the generated text (Raffel et al. 2020). The Better & Faster LLMs apply multi-token prediction to LLMs, allowing them to directly predict multiple tokens in a single inference step, thereby reducing the number of inference iterations and associated computational costs (Gloeckle et al. 2024). However, the number of predicted tokens is fixed, which may inadvertently combine unrelated tokens, potentially disrupting the original semantics of the sentence and consequently decreasing inference accuracy.
Encoder-decoder model: The concept of an encoder is introduced by Yann LeCun in his doctoral dissertation, wherein he posited that an encoder could transform input data into shorter representations without losing the core information, thereby providing a foundational approach to compressing high-dimensional data. The Variational Autoencoder (VAE) model introduces the use of hidden variables between the encoder and decoder to learn the data distribution, laying the groundwork for deep-learning-based encoders and decoders (Kingma and Welling 2013). Sutskever et al. introduces the encoder-decoder framework into the field of deep learning, utilizing recurrent neural networks (RNNs) as both the encoder and decoder for sequence-to-sequence (Seq2Seq) learning, demonstrating exceptional performance in machine translation tasks (Sutskever, Vinyals, and Le 2014). The SegNet applies the encoder-decoder model to the field of image segmentation, significantly enhancing segmentation performance (Badrinarayanan, Kendall, and Cipolla 2017). Subsequently, the encoder-decoder model has achieved great success in various domains including machine translation (Cho et al. 2014; Makin, Moses, and Chang 2020), text-to-speech conversion (Li et al. 2020), computational imaging (Chen et al. 2023; Li et al. 2024a), and beyond, fully demonstrating the capability of the encoder-decoder architecture to efficiently compress and restore data.
By leveraging the advantages of multi-token prediction and the encoder-decoder model, we propose SentenceVAE. SentenceVAE is capable of compressing the information contained within a sentence into a single token, thereby enabling next-sentence prediction and significantly accelerating the inference speed of LLMs. Additionally, This approach divides the text into sentences without compromising the original semantics, thereby accelerating the inference speed of LLMs while ensuring inference accuracy.
Method
In this section, we introduce the SentenceVAE framework. Embedding this framework into a LLM enables the model to implement a sentence-by-sentence prediction method, significantly accelerating its inference speed. The inference process of the proposed method compared to traditional LLMs is illustrated in Fig. 2. It is evident that for the same text, an LLM embedded with SentenceVAE requires only three inference steps to obtain the inference result, whereas traditional LLMs necessitate twenty-free inference steps. Furthermore, the proposed method divides the text into sentences without compromising the original semantics, thereby enhancing the inference speed of LLMs while maintaining inference accuracy.
Sentence Variational Autoencoder (SentenceVAE)
Our SentenceVAE primarily consists of a Sentence Encoder and a Sentence Decoder. The Sentence Encoder encodes multiple word-level tokens from a sentence into a single sentence-level token, while the Sentence Decoder reconstructs this sentence-level token back into the original sequence of word-level tokens. To ensure that the multiple tokens input to the Sentence Encoder come from individual sentences, we propose a specialized Sentence Segmentation Mechanism. Additionally, we introduce a Feature Fusion Mechanism, which enable the encoder to encode variable-length token sequences into a single sentence-level token.
Sentence Segmentation Mechanism: To ensure that all word-level tokens input to the sentence encoder per time come from a single sentence, we use regular matching to split sentences by punctuation marks such as ",", ".", "?", "!" before tokenizing. Let the original string be $S$, and after partitioning, the set $\{s_{i}|1\leq i\leq n\}$ is obtained, satisfying $S=s_{1}s_{2}s_{3}...s_{n}$.
Sentence Encoder: Taking $s=s_{t}(1\leq t\leq n)$ as an example, after tokenization, we obtained a sequence of $L$ token ids $\boldsymbol{D}=[d_{1},d_{2},d_{3},...,d_{L}]$ representing a sentence. This sequence was then passed through an embedding layer, resulting in word-level embedding vectors of $\boldsymbol{E}$, as shown in Eq.1.
$\boldsymbol{E}=(\boldsymbol{e}_{i})_{L\times\text{hidden\_size}}=\textbf{Embed}(\boldsymbol{D})$ (1)
These embeddings were subsequently input into self-attention based encoder blocks, yielding hidden features $\boldsymbol{H}$, as shown in Eq. 2.
$\boldsymbol{H}=(\boldsymbol{h}_{i})_{L\times\text{hidden\_size}}=\textbf{EncoderBlocks}(\boldsymbol{E})$ (2)
To derive the sentence embedding vector, we fused these $L$ features into a single vector $\boldsymbol{\Omega}_{t}\in\mathbb{R}^{\text{hidden\_size}}$. This sentence embedding vector is then fed into the decoder-only LLM, which generates a new sentence embedding vector $\boldsymbol{\Omega}_{t+1}$ at each time step.
Feature Fusion Mechanism: After the encoder blocks generates $\boldsymbol{H}$, we propose a method to fuse them into a single sentence embedding vector $\boldsymbol{\Omega}_{t}$. This method involves accumulating the $L$ vectors and normalizing them using Layer Normalization (Ba, Kiros, and Hinton 2016), as shown in Eq. 3.
$\boldsymbol{\Omega}_{t}=\textbf{LayerNorm}(\sum_{i=1}^{L}{\boldsymbol{h}_{i}})$ (3)
Sentence Decoder: The Sentence Decoder contains a combination of masked self-attention and cross-attention blocks, where the cross-attention mechanism uses the individual sentence embedding vector $\boldsymbol{\Omega}_{t+1}$ as key (K) and value (V).
In the prediction phase, given $\boldsymbol{\Omega}_{t+1}$, initialize the input token $d_{0}=<bos>$, the decoder outputs $d_{1}$, and then input $[d_{0},d_{1}]$ and $\boldsymbol{\Omega}_{t+1}$ into the decoder. Repeat this process until $d_{l^{\prime}+1}=<eos>$, and terminate the iteration, as shown in Eq.4.
$\begin{split}\textbf{init: }&d_{0}=<bos>,l^{\prime}=0\\ \textbf{do: }&\\ &\boldsymbol{D}_{l^{\prime}}=[d_{0},d_{1},d_{2},...,d_{l^{\prime}}]\\ &\boldsymbol{P}_{l^{\prime}}=(\boldsymbol{p}_{i})_{(l^{\prime}+1)\times\text{hidden\_size}}\\ &=\textbf{DecoderBlocks}(\boldsymbol{D}_{l^{\prime}},\boldsymbol{\Omega}_{t+1})\\ &d_{l^{\prime}}=\textbf{argmax}(\boldsymbol{p}_{l^{\prime}})\\ &\boldsymbol{D}_{l^{\prime}+1}=\textbf{concat}(\boldsymbol{D}_{l^{\prime}},d_{l^{\prime}})\\ &l^{\prime}=l^{\prime}+1\\ \textbf{until: }&d_{l^{\prime}}=<eos>\\ \textbf{ret: }&\boldsymbol{D}_{L^{\prime}=l^{\prime}}\end{split}$ (4)
After the iteration is completed, the detokenization process is used to convert the output $\boldsymbol{D}_{L^{\prime}}$ back into a string.
In the training phase, parallel training is adopted, based on masked self-attention mechanism, to ensure that only tokens before $d_{l^{\prime}}$ can be seen when inferring $d_{l^{\prime}+1}$. Given an input sequence of $\boldsymbol{D}_{\text{train}}=[d_{0},d_{1},...,d_{L^{\prime}}]$, model output logits of $\boldsymbol{P}\in\mathbb{R}^{(L^{\prime}+1)\times\text{hidden\_size}}$, and ground truth of $\boldsymbol{D}_{\text{groud\_truth}}=[d_{1},d_{2},...,d_{L^{\prime}},<eos>]$, focal loss (Lin et al. 2017) is taken as the loss function, as shown in Eq. 5.
$\begin{split}\boldsymbol{P}&=\textbf{DecoderBlocks}(\boldsymbol{D}_{\text{train}},\boldsymbol{\Omega_{t+1}})\\ \boldsymbol{P}^{\prime}&=\textbf{softmax}(\boldsymbol{P})\\ \textbf{Loss}&=\sum_{i=1}^{L^{\prime}}-(1-p_{d_{i}})^{\gamma}\log(p_{d_{i}}),\gamma=2\end{split}$ (5)
Sentence-level Large Language Models (SLLMs)
Currently, almost all decoder-only LLMs consist of the following components: an embedding layer $\textbf{EMB}_{\text{llm}}$, $N$ decoder-only blocks $\textbf{DB}_{\text{llm}}$, and a fully connected layer $\textbf{FC}_{\text{llm}}$ for outputting logits. As shown in Eq. 6, given context tokens $\boldsymbol{D}_{\text{context}}=[d_{0},d_{1},...,d_{l^{\prime}}]$, the model generates the next token $d_{l^{\prime}+1}$.
$\begin{split}\boldsymbol{E}_{\text{context}}&=\textbf{EMB}_{\text{llm}}(\boldsymbol{D}_{\text{context}})\\ \boldsymbol{H}_{\text{llm}}&=\textbf{DB}_{\text{llm}}(\boldsymbol{E}_{\text{context}})\\ \boldsymbol{P}_{\text{llm}}&=(p_{\text{llm},i})_{(l^{\prime}+1)\times\text{hidden\_size}}\\ &=\textbf{FC}_{\text{llm}}(\boldsymbol{H}_{\text{llm}})\\ d_{l^{\prime}+1}&=\textbf{argmax}(\boldsymbol{p}_{\text{llm},l^{\prime}})\end{split}$ (6)
In this section, we introduce a unified grafting method to graft SentienceVAE onto the beginning and end of any LLM. This approach transforms the LLMs into Sentence-level Large Language Models (SLLMs), enabling it to operate effectively in the sentence-level embedding space.
Discard Embedding Layer: In a SLLM, the $N$ decoder-only blocks no longer receives word-level tokens directly. Instead, it receives sentence embedding vectors encoded by the sentence encoder. Consequently, the traditional embedding layer in the LLM architecture is removed. Therefore, the model follows Eq. 7.
$\begin{split}\boldsymbol{E}_{\text{sllm}}&=[\boldsymbol{\Omega}_{1},\boldsymbol{\Omega}_{2},...,\boldsymbol{\Omega}_{t}]\\ \boldsymbol{H}_{\text{sllm}}&=(\boldsymbol{h}_{\text{sllm},i})_{t\times\text{hidden\_size}}\\ &=\textbf{DB}_{\text{llm}}(\boldsymbol{E}_{\text{sllm}})\end{split}$ (7)
Termination judgment layer: In traditional LLMs, the final fully connected layer $\textbf{FC}_{\text{llm}}$ typically converts an $\boldsymbol{H}_{\text{llm}}\in\mathbb{R}^{t\times\text{hidden\_size}}$ to a probability vector (logits) $\boldsymbol{P}_{\text{llm}}\in\mathbb{R}^{t\times V}$, where $V$ represents the size of the vocabulary. This allows the model to determine the next token output using sampling algorithms, e.g., Greedy Search, Beam Search (Graves 2012), Random Sampling, Temperature Sampling (Hinton, Vinyals, and Dean 2015), Top-K Sampling (Fan, Lewis, and Dauphin 2018), and Nucleus Sampling (Holtzman et al. 2019). If the current token output is a special token like eos (end-of-sequence), the iteration terminates.
However, in SLLMs, the output of $\textbf{DB}_{\text{llm}}$ is a sentence-level hidden state vector $\boldsymbol{H}_{\text{sllm}}$. Therefore, traditional token generation methods are not applicable. Instead, we use a new fully connected layer $\textbf{FC}_{\text{sllm}}$ called termination judgment layer that converts the $\boldsymbol{H}_{\text{sllm}}$ into a 2-dimensional boolean vector $\boldsymbol{B}_{\text{stop\_flag}}$, as shown in Eq. 8. This vector helps determine whether the current sentence-level hidden state vector signifies the end of a sentence (stop flag) or needs further decoding by the sentence decoder. If the vector indicates an end flag, the iteration terminates. Otherwise, the sentence decoder processes the embedding to generate corresponding tokens, as shown in Eq. 9.
$\begin{split}\boldsymbol{B}_{\text{stop\_flag}}&=(\boldsymbol{b_{i}})_{t\times 2}\\ &=\textbf{FC}_{\text{sllm}}(\boldsymbol{H}_{\text{sllm}})\end{split}$ (8)
$\boldsymbol{\Omega}_{t+1}=\begin{cases}<eos>&,\textbf{argmax}(\boldsymbol{b_{t}})=1\\ \boldsymbol{h}_{\text{sllm},t}&,\textbf{argmax}(\boldsymbol{b_{t}})=0\end{cases}$ (9)
During the training phase, calculate the focal loss for $\boldsymbol{B}_{\text{stop\_flag}}$ as part of the global loss.
Inferencing sentence by sentence: In the SLLM architecture, we use a small model (SentenceVAE) to encode multiple tokens within a sentence into a single sentence embedding vector. This enables the LLM to conduct inference on a sentence-by-sentence basis rather than the traditional token-by-token approach. Experimental results indicate that this approach can significantly reduce the computational cost required for model inference and enhance inference speed.
Table 1: Metrics of SentenceVAEs on validation set.
[…]
SVAE-768-H1 768 1 1.339 3.605
SVAE-768-H2 768 2 1.019 2.588
SVAE-768-H4 768 4 0.5598 1.649
SVAE-1024-H1 1024 1 0.9266 2.406
SVAE-1024-H2 1024 2 0.6610 1.845
SVAE-1024-H4 1024 4 0.3704 1.384
SVAE-2048-H1 2048 1 0.5165 1.622
SVAE-2048-H2 2048 2 0.2845 1.292
SVAE-2048-H4 2048 4 0.1270 1.115
Table 2: Output examples of SentenceVAEs.
ID Input Sentence Toks Output Sentence
SVAE-768-H4 SVAE-1024-H4 SVAE-2048-H4
1 Hello, 3 Hello, Hello, Hello,
[…]
Experiment
To validate our idea, we first trained individual SentienceVAEs using self supervised methods, demonstrating that multiple tokens in a sentence can be compressed into a single sentence vector by an encoder and restored by a decoder. Subsequently, we grafted the encoder and decoder at the beginning and end of the open-source LLMs, proving that with simple modifications, LLMs can be upgraded to SLLMs and work in sentence embedding space, while preserving PPL and improving inference speed. Meanwhile, by observing the loss curve, we found that SLLMs still follows the Scaling Law.
Experimental Setting
The experiments described in this paper were conducted using either a single 4-card RTX 4090 (24G) or a 4-card A100 (40G) (for SLLM-1.3B) setup, with training carried out in a data parallel distribution.
Base Models: Due to memory and time limitations, we only use the 125m, 350m, and 1.3B models of the OPT series (Zhang et al. 2022) as the base LLMs, and demonstrate the feasibility of extending to larger scale models through Scaling Law (Kaplan et al. 2020).
Dataset: We use the English subset (EN/WebText) of the Wanjuan-1.0 dataset (He et al. 2023) as the training dataset. The number of training samples depends on the number of iterations, with SentenceVAE training samples consisting of approximately 153.6M sentences (approximately 1.7B tokens); The training sample size of SLLM is about 6.4M paragraphs (approximately 5.6B tokens). The validation set consists of 1,000 random sentences or paragraphs that are mutually exclusive with the training set.
Hyperparameters: When training SentenceVAE, we assume each sentence has a maximum of 64 tokens, with a batch size of 128/card, a base learning rate of 1e-7/batch/card, and a maximum of 300K iterations.
For training SLLM, we assume each sentence has a maximum of 64 tokens, each paragraph has a maximum of 64 sentences (due to memory limitations), a batch size of 1/card, a base learning rate of 1e-6/batch/card, and a maximum of 1.6M iterations.
All experiments were conducted using the AdamW optimizer (Loshchilov and Hutter 2017) with AMP enabled (Micikevicius et al. 2017), a weight decay coefficient of 0.01, and a maximum L2 norm of 1 for gradient clipping.
The first 5K iterations utilized a linear learning rate schedule, starting at 1/1000 of the base learning rate. After 5K iterations, a cosine annealing learning rate schedule was employed with a period of 20K iterations. Additionally, all experiments implemented the Exponential Moving Average (EMA) strategy.
Metrics: We use perplexity (PPL) as an evaluation metric. The formula for calculating perplexity for output logits $\boldsymbol{P}$ of Sentence Decoder and ground truth $\boldsymbol{D}_{\text{groud\_truth}}$ is as Eq. 10.
$\textbf{PPL}=\exp{(\frac{1}{L^{\prime}}\sum_{i=1}^{L^{\prime}}-\log(p_{d_{i}}))}$ (10)
(a) Test loss curve of SVAEs.
(b) Test loss curve of SLLMs.
Figure 3: Scaling Law of SVAEs and SLLMs.
A series of word-level tokens can be represented by a sentence-level embedding with strong robustness
In this experiment, we trained individual SentienceVAEs using a self-supervised approach. If the Sentence Encoder can compress multiple tokens from the input sentence into a single sentence embedding vector and the Sentence Decoder can use this vector to reconstruct the original token sequence, it would validate the feasibility of our theory.
We trained SentenceVAEs with hidden sizes of 768, 1024, and 2048, corresponding to OPT-125M, OPT-350M, and OPT-1.3B, respectively. For each hidden size, we varied the number of block layers to 1, 2, and 4. We evaluated the models using cross-entropy loss and perplexity (PPL) metrics on the validation set. The experimental results are presented in Table 1.
The evaluation results demonstrate that SentenceVAEs fulfill their intended purpose, confirming the theoretical validity. Table 2 presents several test cases. Among these, Samples 1, 2, 3, 5, 6 are convention samples. To assess the model’s robustness, we also included several out-of-distribution samples, highlighted in blue in the table:
  • Sample 4, 7: Data preprocessing involves segmenting sentences based on punctuation marks, such as "," and ".". Consequently, these samples will be split into multiple inputs during both training and practical application. Directly inputting these samples without preprocessing is considered invalid input for the model.
  • Sample 8: This is the pinyin of a famous Chinese saying, not English, so it is illegal input.
From the output results of the model, we can analyze that:
  • In the convention samples, except for Sample 6, the model accurately restored the sentences. The reason for the error in Sample 6 may be related to the bottleneck of the sinusoidal position encoding mechanism (Vaswani et al. 2017).
  • Although Sample 4 contains the illegal input, the model has accurately restored it with only punctuation marks. Sample 7 still maintains the meaning of the sentence. In Sample 8, as the hidden size of the model increases, the model also accurately restores the illegal input “yao yao ling xian”. These phenomena indicate that the model is robust to illegal inputs.
Table 3: Benchmark test results of OPTs and SLLMs.
Model Total Params Average PPL Mean output throughput (toks/s) Mean GPU memory (KB/token)
[…]
SLLM-125M-H1 214M 26.75 31.68 +18.4% 214.57 652.78 +204.2% 73.15 12.03 -83.6%
SLLM-125M-H2 226M 44.60 +66.7% 539.80 +151.6% 7.08 -90.3%
SLLM-125M-H4 250M 14.32 -46.5% 332.12 +54.8% 10.00 -86.3%
SLLM-350M-H1 429M 25.18 24.84 -1.4% 144.33 481.39 +233.5% 197.59 29.98 -84.8%
SLLM-350M-H2 450M 14.81 -41.2% 442.23 +206.4% 26.78 -86.4%
SLLM-350M-H4 492M 10.17 -59.6% 315.61 +118.7% 17.73 -91.0%
SLLM-1.3B-H1 1.61B 15.95 8.76 -45.1% 119.07 479.71 +302.9% 400.01 57.07 -85.7%
SLLM-1.3B-H2 1.69B 3.84 -75.9% 553.95 +365.2% 55.14 -86.2%
LLMs can work in setence-level embedding space with faster inference speed, more accurate PPL, and longer context, resulting SLLMs
After verifying the feasibility of Sentence level embedding, we grafted the encoder and decoder of SentenceVAE onto the beginning and end of the OPT series models, corresponding to the relationship shown in Table 4.
Table 4: The correspondence relationship between SLLMs, base LLMs, and SentenceVAEs
Model Base Model Grafted SentenceVAE
SLLM-125M-H1 OPT-125M SVAE-768-H1
SLLM-125M-H2 SVAE-768-H2
SLLM-125M-H4 SVAE-768-H4
SLLM-350M-H1 OPT-350M SVAE-1024-H1
SLLM-350M-H2 SVAE-1024-H2
SLLM-350M-H4 SVAE-1024-H4
SLLM-1.3B-H1 OPT-1.3B SVAE-2048-H1
SLLM-1.3B-H2 SVAE-2048-H2
We did not test SLLM-1.3B-H4 due to memory overflow issues encountered during data-parallel distributed training. The benchmark test results of OPTs and SLLMs are shown in Table 3.
Faster inference speed: We assessed the throughput rate of SLLMs using the PyTorch framework (Imambi, Prakash, and Kanagachidambaresan 2021), without incorporating optimization methods such as PagedAttention (Kwon et al. 2023) or acceleration engines like TensorRT-LLM or LMDeploy (Contributors 2023). The throughput rate of the original OPT model served as the baseline for comparison. The experimental results show that SLLMs is 2$\sim$3$\times$ faster on average than LLMs. The larger the number of model parameters, the smaller the proportion of inferencing cost occupied by SentenceVAE, and the more obvious the speed benefit.
More accurate PPL: We observed that the PPL metric for SLLMs was better than that of the baseline OPT models. This improvement is likely due to the SLLM paradigm enabling the model to process natural language at a higher sentence level, leading to enhanced performance. Additionally, the attention computation in the sentence embedding space allows SLLMs to manage shorter relative contexts for texts of the same length, which further contributes to the improved model performance.
Longer context length: The context length that a model can handle is primarily constrained by GPU memory. Under the SLLM framework, multiple original tokens can be compressed into a single token, allowing the model to use less memory when processing contexts of the same length. This effectively increases the context length that the model can handle given the same hardware resources.
SLLMs still follow the Scaling Law
By analyzing the loss curve of SLLM during the training process, we observed that it adheres to the Scaling Law (Kaplan et al. 2020), as shown in Fig. 3. This suggests that the SLLM framework can be effectively extended to larger-scale language models through scaling.
Conclusion & Future Trends
In this section, we summarize our work and suggest several potential directions for future development.
Scaling up SLLMs with better structures
In this work, to rapidly validate our ideas, we utilized the classic Transformer architecture (Vaswani et al. 2017) implemented in the PyTorch framework (Imambi, Prakash, and Kanagachidambaresan 2021). Due to constraints in computing power and time, we limited our validation to models ranging from 125M to 1.3B parameters. However, by verifying the Scaling Law (Kaplan et al. 2020), we inferred the applicability of our approach to larger scale models. Numerous state-of-the-art architectural optimizations can be integrated into the SLLM framework, such as Rotational Position Encoding (RoPE) (Su et al. 2024), which may address the order problem observed in Sample 6. Currently, our models are trained solely on English corpora, but there is potential to add multilingual support in the future. In summary, while the SLLM framework is still in its early stages, it holds significant potential for further optimization.
SLLMs assist in hybrid edge-cloud inference
The SLLM framework is inherently a hybrid architecture combining small and large models. By identifying the optimal balance point, the small model (SentenceVAE) and the large model (LLM) can be effectively distributed between the edge and cloud, balancing computational load and enhancing user experience.
SLLMs assist in embodied intelligence
Currently, embodied intelligence based on the LLM Agent paradigm struggles to interact directly with underlying hardware, relying instead on intermediate components to translate high-level LLM instructions into traditional real-time control logic. The primary limitation is the insufficient rate of “token” generation on the edge side. SLLM, however, can process and generate more “tokens” within the same computational and temporal constraints, offering the potential for embodied intelligent large models to interface directly with underlying hardware.
SLLMs assist in multimodal large model
The “frames” in videos and “trunks” in audio are, to some extent, equivalent to “tokens” in language models. Therefore, the SLLM paradigm can similarly be applied to multimodal large models to increase the processing “frame rate”, enhance user experience, and achieve performance comparable to or even surpassing that of GPT-4o.